# Stronger ViTs With Octic Equivariance

### v2

@@ 3.1 Preliminaries on Octic Equivariance
$b(\rho_{\text{token}}(g)\mathbf{x})=\rho_{\text{token}}(g)b(\mathbf{x}).$ (7)
@@
$b\left(\rho_{\text{token}}(g)\mathbf{x}\rho_{\text{chan}}(g)^{\mathsf{T}}\right)=\rho_{\text{token}}(g)b(\mathbf{x})\rho_{\text{chan}}(g)^{\mathsf{T}}.$ (8)
In terms of an individual token $x\in\mathbb{R}^{D}$ transforming according to $\rho_{\text{chan}}$ we will write this as
$b\left(\rho_{\text{chan}}(g)x\right)=\rho_{\text{chan}}(g)b(x).$ (9)
Designing the components of octic ViT blocks so that they satisfy (8) is the topic of Section 3.2.
3.1.1 Notation
For the convenience of the reader, we collect the most important notation in the paper in this section. We use the bold letter $\mathbf{x}$ for ViT features, which have shape $L\times D$, where $L$ is the number of tokens and $D$ the channel dimension. Typically, $L=N^{2}$, where $N$ is the image height/width in tokens. For an individual $D$-dimensional token we use the letter $x$ which is often acted on by $\rho_{\text{chan}}=\frac{D}{8}\rho_{\text{iso}}$, in which case we can split $x$ into $D/8$-dimensional sub-tokens $x_{\text{A1}},x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E11}},x_{\text{E12}},x_{\text{E21}},x_{\text{E22}}$, where the first four transform according to the irreps $\rho_{\text{A1}},\rho_{\text{A2}},\rho_{\text{B1}},\rho_{\text{B2}}$ respectively while the $2\times\frac{D}{8}$-matrices $(x_{\text{E11}}\quad x_{\text{E12}})^{\mathsf{T}}$ and $(x_{\text{E21}}\quad x_{\text{E22}})^{\mathsf{T}}$ both transform according to $\rho_{\text{E}}$.
@@ 3.2 Octic Layers
LN is layer normalization, MHA multi-head self-attention and MLP a $D\to 4D$ linear map, GELU, and $4D\to D$ linear map. The blocks can subsequently be stacked, as illustrated in Figure 2.
@@ 3.2.1 The Patch Embedding Layer
The first layer in a ViT, following [21], is the patch embedding, short PatchEmbed. In our case, it can be viewed as a mapping from steerable features of type $\rho_{\text{patch}}$ to steerable features of type $\frac{D}{8}\rho_{\text{iso}}$ where $D$ is the embedding dimension of the ViT. It is implemented as a convolution over the input image with kernel size and stride equal to the patch size $P$. The convolution kernels are weight sharing constrained to map to features of the different irreps-types in $\frac{D}{8}\rho_{\text{iso}}$, see Figure 3 for an illustration.
Directly after PatchEmbed, we add a learnable positional encoding $\mathbf{e}\in\mathbb{R}^{L\times D}$ to the features. The positional encoding is not constant over tokens, thereby breaking translation equivariance. To be $\mathrm{D}_{8}$ equivariant $\mathbf{e}$ must satisfy
$\begin{split}&\rho_{\text{token}}(g)\mathbf{x}\rho_{\text{chan}}(g)^{\mathsf{T}}+\mathbf{e}=\rho_{\text{token}}(g)(\mathbf{x}+\mathbf{e})\rho_{\text{chan}}(g)^{\mathsf{T}}\quad\Longleftrightarrow\quad\mathbf{e}=\rho_{\text{token}}(g)\mathbf{e}\rho_{\text{chan}}(g)^{\mathsf{T}}.\end{split}$ (11)
In words, the positional encoding at a specific token position $p$ must be a $\rho_{\text{chan}}(g)$-transformed version of the positional encoding at the token position that is permuted to $p$ by $\rho_{\text{token}}(g)$.
After adding the positional encoding, we append a learnable class token [CLS]$\in\mathbb{R}^{1\times D}$ [19] to the features. To ensure equivariance, we enforce it to be non-zero only in the A1 feature type.
@@ 3.2.2 Linear Layers
Linear layers appear in ViTs both in the MLP block and the MHA block. As mentioned in Section 3.1, equivariant linear layers map between irreps of the same type due to Schur’s lemma. This fact was used to construct efficient reflection-equivariant neural networks by Bökman et al. [8]. Here, we use the same approach for octic equivariance.
To re-iterate, for features of type $\rho_{\text{chan}}=\frac{D}{8}\rho_{\text{iso}}$ we consider each token $x\in\mathbb{R}^{D}$ split into $x_{\text{A1}},x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E}}$ of dimensions $D/8,D/8,D/8,D/8$ and $42\times D/$ 4. Linear layers map each irrep type to itself, meaning that they are parameterised by four $D/8\times D/8$ matrices and one $D/4\times D/4$ matrix, yielding a factor $8$ fewer parameters than a general linear layer.
[…]
@@ 3.2.3 Activation Functions, Layer Norm, Attention and Invariantization
A pointwise activation function $\sigma$ can be applied equivariantly after transforming the features from the Fourier domain (multiples of $\rho_{\text{iso}}$), to the spatial domain (multiples of $\rho_{\text{reg}}$), as discussed in Example 3.3. In the spatial domain $\sigma$ can be applied point-wise as this commutes with the permutation representation $\rho_{\text{reg}}$. In our ViTs, following prior work, we apply the GELU [27] activation function.
We implement (token-wise) equivariant layer normalization by transforming each irrep separately to mean 00 followed by division by the total norm over all the irreps.
@@
To output $\mathrm{D}_{8}$ invariant classification predictions, we map from features of type $\frac{D}{8}\rho_{\text{iso}}$ to features of type $D\rho_{\text{A1}}$ and then use global average pooling over tokens. We ablate different invariantizations to A1-tokens in Appendix C, including linear invariants, triple correlation [43, 29], max filtering [9], generators of the ring of invariant polynomials and canonisation of the signal [28]. We find that a power spectrum invariantization works well and settle on that for the remainder of the experiments.
@@ 3.3 Computational Efficiency
As Transformers scale, the linear layers dominate the execution time [30]. Thus, as ViTs grow, the FLOPs savings will approach those of the linear layer, a reduction of $5.33$ times. We plot the FLOPs savings of a ViT block as the embedding dimension increases in Figure 4(a). The computational benefits of octic ViTs are more pronounced at scale, and we benchmark the throughput of various ViT sizes from the literature and their octic counterparts in Table 1.
Table 1: Compute scaling. We measure the scaling of octic ViTs. The model sizes are taken from Dehghani et al. [17] and we do not train the largest models as part of this work. The numbers ending in “x” describe the improvement over standard ViT statistics of the corresponding octic ViTs.
Model Width Depth MLP Heads throughput FLOPs Peak Mem.
ViT-L 1024 24 4096 16 1.33x 4.58x 2.44x
ViT-H 1280 32 5120 16 1.57x 4.58x 2.88x
@@
As shown in Table 1, savings in FLOPs do not translate one-to-one to improvements in throughput (images per second) in our current implementation. However, it is still the case that the throughput is greatly improved. Our models are pure PyTorch with torch.compile, except for the GELU nonlinearity. We implement a custom Triton [45] kernel that fuses the GELU nonlinearity with the Fourier and inverse Fourier transforms, limiting memory transfers and kernel invocation overhead. While ordinary GELU is pointwise, the new fused Triton kernel is eight points to eight points. For extra efficiency we implement the Fourier transforms by a FFT on $\mathrm{D}_{8}$, described in Appendix A.
Finally, we study the arithmetic intensity of our octic linear layer. The arithmetic intensity measures FLOPs per transferred byte and can be compared with the FLOPs per bandwidth of a given device to obtain a bound on the maximum achievable throughput [2]. The arithmetic intensities are $\frac{2BDF}{P(BD+DF+BF)}$ and $\frac{2BDF/5.33}{P(BD+DF/8+BF)}$ for the standard and octic linear layers, respectively, where $B$ is the batch size in tokens, $D$ is the input dimension, $F$ is the output dimension and $P$ is the precision in bytes. This means that the octic and ordinary layers scale differently. At high scale, not only FLOPs are improved by octic layers but also arithmetic intensity. For instance, for $B=196$ (one image worth of tokens), $P=2$ and $F=4D$ (a typical MLP expansion factor) one can calculate that ordinary linear layers have higher arithmetic intensity up to $D\approx 3200$, whereas octic linear layers have higher arithmetic intensity at larger dimensions. For the experiments in this paper, we are not able to scale to such large dimensions, but still get throughput benefits due to savings in FLOPs, as shown in Table 1.
4 Experiments
In this section, we begin with an ablation study of the number of octic blocks followed by an empirical evaluation of our octic ViTs on both supervised (DeiT III) and self-supervised (DINOv2) tasks. We will release code and weights on GitHub.
4.1 Training algorithms and data
DeiT III
[46] (Data Efficient Image Transformers) is a popular and highly tuned supervised training recipe for classification with ViTs on ImageNet. We follow the official implementation.
DINOv2
[37] is a state-of-the-art self-supervised method to extract visual features. It combines iBOT [60] and DINO [11] objectives at scale. We follow the official implementation.
All of our models are trained on ImageNet-1K [18, 42, 40] with recipes that match the original implementations. In Appendix D, we provide further details on the experiments.
4.2 Ablation of number of octic blocks
We experiment with incorporating non-equivariant blocks [52, 8]. We include networks where the first $k$ of the ViT blocks are octic and the remaining $l-k$ are standard blocks. When we break equivariance in the transition from the equivariant to standard blocks we call the network hybrid (denoted $\mathcal{H}_{p}$) and when we map to $D\rho_{\text{A1}}$-steerable features in the transition we refer to the network as early invariant (denoted $\mathcal{I}_{p}$). We denote networks with only equivariant layers (mapping to invariant features after the last layer) by $\mathrm{D}_{p}$, where $p$ refers to order of the dihedral group. Figure 4(b) ablates different values of $k$ for ViT-L ($l=24$). We find that $k=\frac{l}{2}$ strikes a good balance between computational efficiency and representational power and use that for the remainder of the experiments. Note that $\mathcal{I}_{8}$ performs worse for small $k$.
@@ 4.3 DeiT III
Table 2: DeiT III evaluation. We measure the Top-1 classification accuracy on ImageNet-1K for different model sizes. We further highlight the computational complexity of each model.
Model params throughput FLOPs Peak Mem. Top-1 OOD Rot.
(×$\times 10^{6}$) (im/s) (×$\times 10^{9}$) (MB) Acc. $\uparrow$ $\Delta$Acc.$\uparrow$
ViT-H/14 632.1 526 167.8 3285 84.6 -12.6
[…]
ViT-L/16 304.4 1317 61.9 1557 84.2 -13.4
[…]
Table 3: DINOv2 evaluation. We evaluate the frozen DINOv2 features by classification accuracy on ImageNet-1K (IN1K) and segmentation mIoU on ADE20K [58, 59] and VOC2012 [23].
@@
(×$\times 10^{9}$) linear $k$-NN linear $k$-NN linear $k$-NN
@@
[…]
@@
[…]
@@ 4.4 DINOv2
[…]
@@ 5 Limitations
In this work, we limit our scope to an extensive study of the $\mathrm{D}_{8}$ group and leave larger dihedral groups for future work. We do not perform hyperparameter tuning and instead replicate the baseline training recipes. Furthermore, we do not conduct an extensive ablation study of the share of features per irrep. We aim to explore this in future work.
We do not realize the full throughput potential of our octic layers, as illustrated by lower throughput gains than FLOPs savings. Our results show promise in continuing work in this direction. Due to computational constraints, we do not replicate the full training of state-of-the-art models such as DINOv2 ViT-G/14. We find that the benefits of octic equivariance increase with model size and expect this trend to persist at even larger scales, which we leave to future work to validate.

### v4

@@ 3.2 Octic ViTs
$b(\mathbf{x}\rho_{\text{token}}(g)^{\mathsf{T}})=b(\mathbf{x})\rho_{\text{token}}(g)^{\mathsf{T}}.$ (7)
@@
$b\left(\rho_{\text{chan}}(g)\mathbf{x}\rho_{\text{token}}(g)^{\mathsf{T}}\right)=\rho_{\text{chan}}(g)b(\mathbf{x})\rho_{\text{token}}(g)^{\mathsf{T}}.$ (8)
Designing the components of octic ViT blocks so that they satisfy equation 8 is the topic of Section 3.3.
Figure 3: Architecture. Patches are first extracted from an image using specialized octic filters and the resulting features are processed by $k$ octic ViT blocks. The final embeddings can be fed to lkl-k standard Transformer blocks (as demonstrated by our $\mathcal{H}_{8}$ and $\mathcal{I}_{8}$ ViTs). When $k=l$, we denote $\mathcal{I}_{8}(\text{ViT})$ by $\mathrm{D}_{8}(\text{ViT})$, which hence only uses octic ViT blocks before a final invariantization.
@@ 3.3 Octic Layers
LN is layer normalization, MHA multi-head self-attention and MLP a $C\to 4C$ linear map, GELU, and $4C\to C$ linear map. The blocks can subsequently be stacked, as illustrated in Figure 3.
@@ 3.3.1 The Patch Embedding Layer
The first layer in a ViT, following (Dosovitskiy et al., 2021), is the patch embedding, short PatchEmbed. In our case, it can be viewed as a mapping from steerable features of type $\rho_{\text{patch}}$ to steerable features of type $\frac{C}{8}\rho_{\text{iso}}$ where $C$ is the embedding dimension of the ViT. It is implemented as a convolution over the input image with kernel size and stride equal to the patch size $P$. The convolution kernels are weight sharing constrained to map to features of the different irreps-types in $\frac{C}{8}\rho_{\text{iso}}$, see Appendices C and C.1 for visualizations of the kernels.
Directly after PatchEmbed, we add a learnable positional encoding $\mathbf{e}\in\mathbb{R}^{C\times L}$ to the features. The positional encoding is not constant over tokens, thereby breaking translation equivariance. To be $\mathrm{D}_{8}$ equivariant $\mathbf{e}$ must satisfy
$\begin{split}&\rho_{\text{chan}}(g)\mathbf{x}\rho_{\text{token}}(g)^{\mathsf{T}}+\mathbf{e}=\rho_{\text{chan}}(g)(\mathbf{x}+\mathbf{e})\rho_{\text{token}}(g)^{\mathsf{T}}\quad\Longleftrightarrow\quad\mathbf{e}=\rho_{\text{chan}}(g)\mathbf{e}\rho_{\text{token}}(g)^{\mathsf{T}}.\end{split}$ (10)
[…]
@@ 3.3.2 Linear Layers
Linear layers appear in ViTs both in the MLP block and the MHA block. As mentioned in Section 3.1 and illustrated in Figure 2, equivariant linear layers map between irreps of the same type due to Schur’s lemma. This fact was used to construct efficient reflection-equivariant neural networks by Bökman et al. (2025). Here, we use the same approach for octic equivariance.
To re-iterate, for features of type $\rho_{\text{chan}}=\frac{C}{8}\rho_{\text{iso}}$ we consider each token $x\in\mathbb{R}^{C}$ split into $x_{\text{A1}},x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E}}$ of dimensions $C/8,C/8,C/8,C/8$ and $2\times C/4$. Linear layers map each irrep type to itself, meaning that they are parameterised by four $C/8\times C/8$ matrices and one $C/4\times C/4$ matrix, yielding a factor $8$ fewer parameters than a general linear layer.
In terms of FLOPs needed to compute the linear layer, $x_{i}\mapsto W_{i}x_{i}$ requires $C/8\cdot C/8=C^{2}/64$ FLOPs for $i\in\{\text{A1, A2, B1, B2}\}$ while due to the “weight-sharing” over the two dimensions in $\rho_{\text{E}}$, it requires $2\cdot C/4\cdot C/4=C^{2}/8$ FLOPs for $i=\text{E}$. In total we therefore get $16/3\approx 5.33$ times fewer FLOPs than the $C^{2}$ required for an ordinary linear layer.
@@ 3.3.3 Activation Functions, Layer Norm, Attention and Invariantization
A pointwise activation function $\sigma$ can be applied equivariantly after transforming the features from the Fourier domain (multiples of $\rho_{\text{iso}}$), to the spatial domain (multiples of $\rho_{\text{reg}}$), as discussed in Example 3.3. In the spatial domain $\sigma$ can be applied point-wise as this commutes with the permutation representation $\rho_{\text{reg}}$. In our ViTs, following prior work, we apply the GELU activation function.
We implement (token-wise) equivariant layer normalization by transforming each irrep separately to mean 0 followed by division by the total norm over all the irreps.
@@
To output $\mathrm{D}_{8}$ invariant classification predictions, we map from features of type $\frac{C}{8}\rho_{\text{iso}}$ to features of type $C\rho_{\text{A1}}$ and then extract the [CLS] token. We ablate different invariantizations to A1-tokens in Appendix D, including linear invariants, triple correlation (Kakarala, 2012; Sanborn & Miolane, 2023), max filtering (Cahill et al., 2024), generators of the ring of invariant polynomials and canonisation of the signal (Kaba et al., 2023). We find that a power spectrum invariantization works well and settle on that for the remainder of the experiments.
@@ 3.4 Computational Efficiency
As Transformers scale, the linear layers dominate the execution time (Kaplan et al., 2020). Thus, as ViTs grow, the FLOPs savings will approach those of the linear layer, a reduction of $5.33$ times. We plot the FLOPs savings of a ViT block as the embedding dimension increases in Figure 4(a). The computational benefits of octic ViTs are more pronounced at scale, and we benchmark the throughput of various ViT sizes from the literature and their octic counterparts in Table 1. We further compare the arithmetic intensity of standard and octic linear layers in Appendix B.
Table 1: Compute scaling. We measure the scaling of octic ViTs. The model sizes are taken from Dehghani et al. (2023) and we do not train the largest models as part of this work. The numbers ending in “x” describe the improvement over standard ViT statistics of the corresponding octic ViTs.
Model Width Depth MLP dim Heads throughput FLOPs Peak Mem.
ViT-L 1024 24 4096 16 1.32x 4.58x 2.44x
ViT-H 1280 32 5120 16 1.47x 4.58x 2.88x
@@
As shown in Table 1, savings in FLOPs do not translate one-to-one to improvements in throughput (images per second) in our current implementation. However, it is still the case that the throughput is greatly improved. Our models are pure PyTorch with torch.compile, except for the GELU nonlinearity. We implement a custom Triton (Tillet et al., 2019) kernel that fuses the GELU nonlinearity with the Fourier and inverse Fourier transforms, limiting memory transfers and kernel invocation overhead. While ordinary GELU is pointwise, the new fused Triton kernel is eight to eight points. For extra efficiency, we implement the Fourier transforms by a FFT on $\mathrm{D}_{8}$, described in Appendix A.
@@
4 Experiments
In this section, we evaluate our octic ViTs on supervised (DeiT III) and self-supervised (DINOv2) training recipes and perform ablations. DeiT III (Touvron et al., 2022) is a popular and highly tuned supervised training recipe for classification and DINOv2 (Oquab et al., 2024) is a state-of-the-art self-supervised method to extract visual features at large scale. All models are trained on ImageNet-1K (Deng et al., 2009; Russakovsky et al., 2015; Recht et al., 2019) following the official implementations. We will release code and pretrained weights on GitHub.
@@ 4.1 DeiT III
We train an array of networks on the supervised task of image classification and compare to the performance reported by Touvron et al. (2022) and Bökman et al. (2025). We find, as illustrated in Table 2, that incorporating octic-equivariant layers provides significant computational savings without sacrificing accuracy. In particular, our $\mathcal{H}_{8}(\text{ViT-H/14})$ model achieves a classification performance of 85.0% compared to the baseline of 84.6% while using only 61% of the FLOPs and matching the performance of the $\mathcal{H}_{2}(\text{ViT-H/14})$ that incorporates flopping ($\mathrm{D}_{2}$) equivariance while being more computational efficient. Similar computational gains are achieved by the invariant model $\mathcal{I}_{8}(\text{ViT-H/14})$, which achieves a classification performance of 84.7%.
Table 2: DeiT III evaluation. We measure the Top-1 classification accuracy on ImageNet-1K for different model sizes. Our networks are marked with $\dagger$.
Model params throughput FLOPs Peak Mem. Top-1 OOD Rot.
@@
ViT-H/14 632.1 569 167.8 3285 84.6 -12.6
$\mathcal{H}_{2}(\text{ViT-H/14})$ 474.2 640 127.4 2734 85.0 -13.4
$\mathcal{I}_{8}(\text{ViT-H/14})^{\dagger}$ 362.3 657 104.0 2249 84.7 0.0
$\mathcal{H}_{8}(\text{ViT-H/14})^{\dagger}$ 355.8 660 102.3 2223 85.0 -13.4
$\mathrm{D}_{2}(\text{ViT-H/14})$ 316.1 739 87.3 2100 84.4 -13.1
$\mathrm{D}_{8}(\text{ViT-H/14})^{\dagger}$ 80.5 836 36.6 1141 82.0 0.0
ViT-L/16 304.4 1421 61.9 1557 84.2 -13.4
$\mathcal{H}_{2}(\text{ViT-L/16})$ 228.3 1610 46.9 1458 84.5 -13.9
$\mathcal{I}_{8}(\text{ViT-L/16})^{\dagger}$ 175.5 1618 38.5 1210 84.0 0.0
$\mathcal{H}_{8}(\text{ViT-L/16})^{\dagger}$ 171.3 1615 37.7 1194 84.5 -13.6
$\mathrm{D}_{2}(\text{ViT-L/16})$ 152.2 1868 32.2 1173 83.4 -14.3
$\mathrm{D}_{8}(\text{ViT-L/16})^{\dagger}$ 39.2 1877 13.5 689 79.5 0.0
Table 3: DINOv2 evaluation. We evaluate the frozen DINOv2 features by classification accuracy on ImageNet-1K (IN1K) and segmentation mIoU on ADE20K (Zhou et al., 2017b; 2019) and VOC2012 (Everingham et al., 2010). Our networks are marked with $\dagger$.
@@
($\times 10^{9}$) linear $k$-NN linear $k$-NN linear $k$-NN
@@
$\mathcal{H}_{2}(\text{ViT-H/16})$ 97.0 81.9 80.9 34.8 30.8 70.7 61.2
$\mathcal{I}_{8}(\text{ViT-H/16})^{\dagger}$ 77.7 81.9 80.9 33.9 29.2 70.6 61.2
$\mathcal{H}_{8}(\text{ViT-H/16})^{\dagger}$ 77.5 82.2 81.4 35.1 31.1 70.8 61.7
@@
$\mathcal{H}_{2}(\text{ViT-L/16})$ 46.9 81.3 80.7 33.4 29.0 69.3 58.0
$\mathcal{I}_{8}(\text{ViT-L/16})^{\dagger}$ 38.5 81.2 80.4 32.6 28.0 70.0 59.6
$\mathcal{H}_{8}(\text{ViT-L/16})^{\dagger}$ 37.7 81.3 80.8 33.6 29.4 69.1 59.5
$\mathrm{D}_{8}(\text{ViT-L/16})^{\dagger}$ 13.5 72.2 71.0 23.7 17.6 53.2 42.5
@@ 4.2 DINOv2
As another pretraining task, we consider the DINOv2 recipe and train our own baselines. Results are summarized in Table 3. We find that incorporating octic-equivariant layers maintains or improves downstream classification and segmentation performance while saving FLOPs. In particular, our invariant network $\mathcal{I}_{8}(\text{ViT-H/16})$ matches the downstream performance of the baseline while using only $61\%$ of the FLOPs and $\mathcal{H}_{8}(\text{ViT-H/16})$ slightly improves performance with similar savings.
In Appendix E.3, we further investigate the performance of our invariant model and evaluate it on white blood cell classification, a task that lacks a canonical orientation. The invariant model $\mathcal{I}_{8}(\text{ViT-L/16})$ outperforms the baseline on most evaluated metrics.
4.3 Ablations
Impact of equivariance.
We study the effect of equivariance by replacing the kernel constrained equivariant patch embed by an arbitrarily linear mapping while keeping the rest of the architecture the same as for $\mathcal{H}_{8}$. In principle, the model can learn to be equivariant. We evaluate ViT-B using the DeiT III recipe and achieve an accuracy of $82.4$ compared to $83.0$ for $\mathcal{H}_{8}$(ViT-B). The results suggest that equivariance yields higher accuracy than arbitrary mappings with the same block-diagonal structure (in addition to providing steerable features that can be useful in downstream applications).
Number of octic blocks.
We experiment with incorporating non-equivariant blocks (Weiler & Cesa, 2019). We include networks where the first $k$ of the ViT blocks are octic and the remaining lkl-k are standard blocks. Figure 4(b) ablates different values of $k$ for ViT-L ($l=24$). We find that $k=\frac{l}{2}$ strikes a good balance between computational efficiency and representational power and use this throughout the paper. Note that $\mathcal{I}_{8}$ performs worse for small $k$ due to early invariantization.
@@ 5 Limitations
In this work, we limit our scope to an extensive study of the $\mathrm{D}_{8}$ group and leave larger dihedral groups for future work. We follow baseline training recipes without hyperparameter tuning, and we do not conduct an extensive ablation study of the share of features per irrep or scale the size beyond ViT-H. We aim to explore this in future work. Moreover, we do not realize the full throughput potential of our octic layers, as illustrated by lower throughput gains than FLOPs savings. Our results show promise in continuing work in this direction.